OpenStudy (anonymous):

Divide. 2x^5 -6x^2+4x^4y over 2x

OpenStudy (lgbasallote):

factor out 2x from the numerator...so the denom can cancel out...

OpenStudy (saifoo.khan):

Simplify\[\frac{2x^5}{2x} - \frac{6x^2}{2x} + \frac{4x^4y}{2x} \]Now cancel out.

OpenStudy (espex):

Factor out 2x from each term and cancel the denominator.

OpenStudy (espex):

haha :)

OpenStudy (saifoo.khan):

;)

OpenStudy (lgbasallote):

^_^